Solve the following measurements:


  1. On a certain map, the straight-line distance between Pas ans Winnipeg is 4.1cm. If the scale of the map is 1cm=120cm, how far apart are the Pas and Winnipeg, to the nearest kilometer?
  2. On a scale drawing of a house, the scale is 1cm:4cm. If the actual house is 30cm, what is the length of the line in the drawing that represents the width of the house?




1
Expert's answer
2021-08-08T17:34:14-0400

(1)

the straight-line distance between Pas and Winnipeg is 4.1cm.

"on\space a \space scale\space 1cm=120cm\\so \space4.1cm=120\times4.1\\=492cm"

now

"1cm=0.00001km\\492cm=0.00001\times 492\space km\\=0.00492\space km"

(2)

according to given condition

"\frac{length\space on\space paper}{actual\space length}=\frac{1}{4}"

"\frac{length \space of \space house \space in \space drawing}{actual\space house \space length}=\frac{x}{30}"

so now

"\frac{1}{4}=\frac{x}{30}\\4x=30\\x=\frac{30}{4}\\x=7.5\space cm"